OpenStudy (anonymous):

What is the value of 10C7?

OpenStudy (anonymous):

hi

OpenStudy (goldphenoix):

The value of 10C7 is 120.

OpenStudy (anonymous):

I knew it, umm can you teach me how to do this problem?

OpenStudy (anonymous):

it's 6*5*4=120

OpenStudy (anonymous):

thx

OpenStudy (goldphenoix):

10C7 is the number of ways of selecting 7 articles from a total of 10, and is calculated from \[\frac{ 10! }{ 7!(10-7)! }

OpenStudy (goldphenoix):

\[\frac{ 10! }{ 7!(10-7!) }\]

OpenStudy (anonymous):

I never thought of it that way, thanks

OpenStudy (anonymous):

oh ok I figured it all out